OpenStudy (anonymous):

Find next three terms? 1, 2, 4, 7, 8, 10, 25, 26, 28, 79, ___ , ___ , ___

OpenStudy (mathstudent55):

This works with groups of three numbers. Once you know the first number of each group, the next two numbers are obtained by simply adding 1 to the first number then adding 2 to the second number. Start with 1, then add 1, then add 2; that gives you 1, 2, 4 To start the second group, notice the first number is 7. 7 is 6 more than 1. Now add 6 to the first number, 6 + 1 = 7, now add 1 then add 2: 7, 8, 10 To start the third group, 25 is 18 more than 7. 18 = 3*6, and 6 was the previous increment. Now add 18 = 3*6 to the first number of second group: 25. Add 1, then add 2: 25, 26, 28 Now add 3*18=54 to the first number of the third group: 25+54 = 79; add 1 then add 2: 79, 80, 82 Now add 3*54 to the first number of the third group: 79 + 3*54 = 241 1, 2, 4, 7, 8, 10, 25, 26, 28, 79, 80, 82, 241

OpenStudy (anonymous):

thanks! is there a way to find the nth term?